# Hallucination Detection in Large Language Models with Metamorphic Relations

### v1

@@ 6.3.1. Mutation Generation Accuracy
The effectiveness of our approach is heavily dependent on the accuracy of the generated mutations. If these mutations do not accurately represent plausible variations of the original responses, subsequent hallucination detection may yield unreliable results. Furthermore, LLMs may introduce minor hallucinations during this process, with double negations in antonym mutation generation being a notable example (42; 5). Indeed, we sometimes encountered a double negative in the antonym mutation-generating process (an example is detailed in Fig 11). This could result in false positives or negatives, ultimately affecting the validity of our findings. Conversely, utilizing more accurate mutation generation methods could significantly enhance the performance metrics of MetaQA.
@@ 7. Related Work
Detecting hallucination in LLMs is imperative for assuring the reliability and trustworthiness of the generated content. A direct strategy involves comparing the model-generated output against reliable knowledge sources (15; 7; 16; 6; 14; 18). Such methods, however, require access to external databases and can have considerable inference costs.
@@
To address this issue in zero-resource settings, several methods have been devised that eliminate the need for retrieval. The fundamental premise behind these strategies is that LLM hallucinations are inherently tied to the model’s uncertainty. The internal states of LLMs can serve as informative indicators of their uncertainty, often manifested through metrics like token probability or entropy (43; 49). When working with open-source LLMs, we can assess the likelihood of hallucination by examining token-level information, such as confidence levels.
@@
However, uncertainty measures require access to token-level probability distributions, which may not be available for models that only provide API access to users, such as ChatGPT (2). Given this constraint, drawing inspiration from legal cross-examination practices, the LMvLM approach was introduced by Cohen et al. (11). This strategy employs an ‘examiner’ LM to question an ‘examinee’ LM, aiming to unveil inconsistencies in claims during multi-turn interactions.
@@
Beyond adopting a multi-agent perspective by incorporating additional LLMs, assessing uncertainty from the self-consistency of a single LLM is often more practical. Additionally, some research has demonstrated the feasibility of this approach. For instance, (4; 46; 21) detect hallucinations through natural language prompts. Moreover, the method proposed by (26) uses logic programming techniques, which are similar to metamorphic relations. However, their approach involves detecting hallucinations under the assumption that the facts are known in advance.
@@
Hallucination evaluation datasets and benchmarks are designed to assess the propensity of LLMs to produce hallucinations, focusing on identifying factual inaccuracies and measuring deviations from the original context. These benchmarks primarily evaluate the factuality of LLM-generated content, often using a question-answering format to emphasize response accuracy. Key benchmark datasets include: TruthfulQA evaluates whether language models generate truthful answers using an adversarial approach to uncover misleading responses from training data. HaluEval (25)is a large-scale hallucination evaluation benchmark for LLMs, featuring a comprehensive collection of generated and human-annotated hallucinated samples. It samples 10K instances from the training sets of HotpotQA, OpenDialKG (33), and CNN/DailyMail (39), targeting question-answering and text summarization tasks. FreshQA addresses hallucinations arising from outdated knowledge, evaluating factuality with 600 hand-crafted questions. It assesses LLMs’ ability to handle fast-changing knowledge and identify questions with false premises. REALTIMEQA (22) emphasizes validating LLMs’ factuality in relation to current world knowledge. It provides real-time open-domain multiple-choice questions from recent news articles across various topics and offers evaluations using accuracy, exact matching, and token-based F1 metrics.

### v2

@@ 6.3.1. Mutation Generation Accuracy
The effectiveness of our approach is heavily dependent on the accuracy of the generated mutations. If these mutations do not accurately represent plausible variations of the original responses, subsequent hallucination detection may yield unreliable results. Furthermore, LLMs may introduce minor hallucinations during this process, with double negations in antonym mutation generation being a notable example (Varshney et al., 2024; Asher and Bhar, 2024). Indeed, we sometimes encountered a double negative in the antonym mutation-generating process (an example is detailed in Fig 11). This could result in false positives or negatives, ultimately affecting the validity of our findings. Conversely, utilizing more accurate mutation generation methods could significantly enhance the performance metrics of MetaQA.
@@ 7. Related Work
Detecting hallucination in LLMs is imperative for assuring the reliability and trustworthiness of the generated content. A direct strategy involves comparing the model-generated output against reliable knowledge sources (Guo et al., 2022; Augenstein et al., 2019; Hanselowski et al., 2019; Atanasova et al., 2020; Galitsky, 2023; Huo et al., 2023). Such methods, however, require access to external databases and can have considerable inference costs.
@@
To address this issue in zero-resource settings, several methods have been devised that eliminate the need for retrieval. The fundamental premise behind these strategies is that LLM hallucinations are inherently tied to the model’s uncertainty. The internal states of LLMs can serve as informative indicators of their uncertainty, often manifested through metrics like token probability or entropy (Varshney et al., 2023; Yao et al., 2023). When working with open-source LLMs, we can assess the likelihood of hallucination by examining token-level information, such as confidence levels.
@@
However, uncertainty measures require access to token-level probability distributions, which may not be available for models that only provide API access to users, such as ChatGPT (cha, 2022). Given this constraint, drawing inspiration from legal cross-examination practices, the LMvLM approach was introduced by Cohen et al. (Cohen et al., 2023). This strategy employs an ‘examiner’ LM to question an ‘examinee’ LM, aiming to unveil inconsistencies in claims during multi-turn interactions.
@@
Beyond adopting a multi-agent perspective by incorporating additional LLMs, assessing uncertainty from the self-consistency of a single LLM is often more practical. Additionally, some research has demonstrated the feasibility of this approach. For instance, (Agrawal et al., 2023; Xiong et al., 2023; Kadavath et al., 2022) detect hallucinations through natural language prompts. Moreover, the method proposed by (Li et al., 2024) uses logic programming techniques, which are similar to metamorphic relations. However, their approach involves detecting hallucinations under the assumption that the facts are known in advance.
@@
Hallucination evaluation datasets and benchmarks are designed to assess the propensity of LLMs to produce hallucinations, focusing on identifying factual inaccuracies and measuring deviations from the original context. These benchmarks primarily evaluate the factuality of LLM-generated content, often using a question-answering format to emphasize response accuracy. Key benchmark datasets include: TruthfulQA evaluates whether language models generate truthful answers using an adversarial approach to uncover misleading responses from training data. HaluEval (Li et al., 2023)is a large-scale hallucination evaluation benchmark for LLMs, featuring a comprehensive collection of generated and human-annotated hallucinated samples. It samples 10K instances from the training sets of HotpotQA, OpenDialKG (Moon et al., 2019), and CNN/DailyMail (See et al., 2017), targeting question-answering and text summarization tasks. FreshQA addresses hallucinations arising from outdated knowledge, evaluating factuality with 600 hand-crafted questions. It assesses LLMs’ ability to handle fast-changing knowledge and identify questions with false premises. REALTIMEQA (Kasai et al., 2024) emphasizes validating LLMs’ factuality in relation to current world knowledge. It provides real-time open-domain multiple-choice questions from recent news articles across various topics and offers evaluations using accuracy, exact matching, and token-based F1 metrics.